lenab:

An artist creates a sculpture out of one ton of salt. If the sculpture loses 3.5% of its mass each year to erosion, after how long will the statue weigh less than 1,000 pounds? using logarithms

justjm:

let y be a function of x, where y is the weight in pounds and x is the time in years. We can depict the scenario through this relation: \(y=(1-0.035)^t\) \(y=0.965^t\) You have to find how long until the statue weighs less than 1000 pounds. \(1000=0.965^t\) Solve by taking the natural log of each side and then using your log rules, then use your calculator.